OpenStudy (poofypenguin):

Simplify the expression: tan [2arccos(x/4)]

OpenStudy (anonymous):

Convert arc cos to arctan..

OpenStudy (poofypenguin):

how would i go about doing that? I've never actually seen inverse trig functions before... i'm not really sure how they work..

OpenStudy (anonymous):

Let ; \[\cos^{-1}(\frac{x}{4}) = y \implies \cos(y) = \frac{x}{4}\] Find tan(y) first..

OpenStudy (anonymous):

Then you should learn what are inverse trigonometric functions and you should remember the formulas related to Inverse trigonometric functions..

OpenStudy (poofypenguin):

Would you recommend any resources? My textbooks have nothing on this...and i haven't seen them in high school before...

OpenStudy (poofypenguin):

I do understand what you did in your previous post, but how would you find tan(y)?

OpenStudy (anonymous):

tan [2arccos(x/4)] Let's see, let y=arccos(x/4) so this means cosy = x/4... So, tan [2arccos(x/4)] = tan [2y] = \(\Large \frac{2tany}{1-tan^2y} \)... Now, look at the triangle in the drawing: